# CRAFT: Calibrated Reasoning with Answer-Faithful Traces via Reinforcement Learning for Multi-Hop Question Answering

### v1

@@ C.3 Output Traces Format Analysis
Table 11 presents a comprehensive analysis of format compliance across all CRAFT variants and model scales.
@@
GRPO dramatically improves format compliance across all scales. Base models struggle with structured output, particularly at smaller scales: 0.5B and 1.5B base models achieve near-zero format accuracy on complex templates (CRAFT–CRAFT), while even 3B models only reach 6.4%–50.3% on average. In stark contrast, GRPO training yields massive improvements: 0.5B models gain +92.5% to +99.7% on CRAFT–CRAFT templates, 1.5B models improve by +60.4% to +99.5%, and 3B models see +31.9% to +95.1% increases. Even 7B base models, despite already achieving 90.9%–99.5% on complex templates (CRAFT–CRAFT), benefit from GRPO with +0.5% to +8.0% gains, reaching near-perfect compliance (97.2%–100%). This demonstrates that reward successfully teaches models to follow structured output requirements across all capacity levels.
@@
Format complexity exhibits scale-dependent learnability. For simpler templates (CRAFT, CRAFT), smaller models show divergent patterns: 0.5B improves dramatically on CRAFT (+98.4%), but exhibits a puzzling anomaly on CRAFT, where GRPO models (0.1%) underperform base (9.0%) by -4.9% to -16.4% across datasets. This suggests the model defaults to more complex formats learned during multi-template training. In contrast, larger models master all formats: 1.5B achieves +80.7% to +99.3% gains, 3B reaches +58.4% to +91.6% improvements, and 7B models show the most dramatic gains on simple templates (+38.2% on CRAFT, +84.6% on CRAFT), indicating that base 7B models paradoxically struggle more with minimal-structure outputs than with complex multi-field templates.
Format Error Analysis. Table 12 presents representative error patterns across all CRAFT variants. Base models predominantly fail due to missing closing tags (60%–85% of errors) and response truncation (30%–40% for complex templates), while GRPO dramatically reduces errors, with CRAFT and CRAFT achieving perfect compliance on 7B models.
@@
<answer>1958 Award missing </answer> Truncation (30%): ...grandmother of Henry St John ( response cut mid-sentence |
@@
<answer> no content, missing </answer> |
| CRAFT |
@@
<answer>[Edwin Schlossberg] missing </answer> Tag mismatch (20%): <reason>...Michael Rubbo </answer> wrong closing tag None bare text |
@@
| CRAFT |
Missing <answer> block (45%):
<reason>...answer is American. </reason> entire <answer>...</answer> missing Truncation (40%): ...cannot determine based on response cut off |
Missing/malformed tags (3%):
<answer>Infanta Maria missing </answer> <answer>The Woman in Fifth </ truncated closing tag |
| CRAFT |
@@
Extra text after tags (0.2%):
<answer>No</answer> (Note: The reasoning...) parenthetical note not allowed |
| CRAFT |
@@
A Bug’s Life no XML tags [Beaumont, Texas] bracket format [No] should be <answer>No</answer> |
@@
C.4 Human Annotation Validation
| Aspect | Both | Both | LLM | LLM | Agr. | |
| Pass | Fail | Human | Human | |||
| Plan-Reason | 388 | 83 | 18 | 11 | 94.2% | 0.81 |
| Evidence Ground. | 380 | 80 | 23 | 17 | 92.0% | 0.75 |
| Answer Deriv. | 400 | 80 | 10 | 10 | 96.0% | 0.86 |
| Gold Doc Citation | 370 | 80 | 30 | 20 | 90.0% | 0.70 |
| Overall | 385 | 80 | 20 | 15 | 93.0% | 0.78 |
To validate the reliability of our LLM-based faithfulness metric, we conducted a human evaluation study following established protocols Adlakha et al. [2024]; Zheng et al. [2023]. Two expert annotators independently assessed 500 randomly sampled MuSiQue reasoning traces from CRAFT with CRAFT template, using the same four-aspect criteria as our LLM judge (Algorithm 1). We measure inter-annotator reliability using Cohen’s Kappa Artstein [2017]:
where is the observed agreement and is the expected agreement by chance. Here denotes both-pass cases, denotes both-fail cases, (LLM pass, human fail) and (LLM fail, human pass) are disagreements, and is the total sample size. Table 13 presents the detailed per-aspect results. The results demonstrate consistently high agreement across all four aspects, with the overall faithfulness achieving 93.0% agreement and between automatic and human evaluation, indicating substantial agreement Zheng et al. [2023]; Bavaresco et al. [2024]. Answer Derivation shows the highest agreement (), as the correctness of final answers is relatively unambiguous. Gold Document Citation has lower agreement (), reflecting occasional ambiguity in determining whether cited documents are strictly necessary. Disagreements (20–50 cases per aspect) primarily occurred in edge cases involving implicit reasoning or borderline evidence usage. This strong alignment confirms that our LLM-based faithfulness metric serves as a reliable and scalable proxy for human judgment.

### v2

@@ B.4 Format Compliance Analysis
Expanding on the format compliance summary in §4.4, Table A presents a comprehensive analysis across all CRAFT variants, model scales, and individual datasets.
@@
GRPO substantially improves format compliance across all scales. Base models struggle with structured output, particularly at smaller scales: 0.5B and 1.5B base models achieve near-zero format accuracy on complex templates, while even 3B models only reach 6.4%–50.3% on average. GRPO training yields massive improvements: 0.5B models gain +92.5% to +99.7% on CRAFT–CRAFT templates, 1.5B models improve by +60.4% to +99.5%, and 3B models see +31.9% to +95.1% increases. Even 7B base models benefit from GRPO with +0.5% to +8.0% gains, reaching near-perfect compliance (97.2%–100%). This demonstrates that successfully teaches models to follow structured output requirements across all capacity levels.
@@
Smaller models struggle with complex templates but benefit most from GRPO training. For simpler templates (CRAFT, CRAFT), smaller models show divergent patterns: 0.5B improves dramatically on CRAFT (+98.4%), but exhibits a notable failure on CRAFT, where GRPO format compliance drops from 9.0% to 0.1%. We attribute this to a reward signal collapse: CRAFT uses only and , but the 0.5B model lacks sufficient capacity to produce correct answers without reasoning structure, yielding near-zero throughout training. With as the only effective reward, the model receives no learning signal to distinguish between the minimal <answer> format and the more complex XML structures encountered during exploration, leading to format drift rather than convergence. This failure is consistent with the complete accuracy collapse of CRAFT on CRAFT (Table A). In contrast, larger models master all formats: 7B models show the most dramatic gains on simple templates (+38.2% on CRAFT, +84.6% on CRAFT), indicating that base 7B models paradoxically struggle more with minimal-structure outputs than with complex multi-field templates.
@@
<answer>1958 Award missing </answer> Truncation (30%): response cut mid-sentence |
@@
<answer> no content |
[…]
@@
<answer>[Edwin Schlossberg] missing </answer> Tag mismatch (20%): </answer> used for </reason> |
@@
[…]
Missing answer block (45%):
entire <answer>...</answer> absent Truncation (40%): response cut off |
Malformed tags (3%):
truncated closing tag |
[…]
@@
Extra text (0.2%):
parenthetical note after tags |
[…]
@@
bare text without XML tags e.g., A Bug’s Life |
@@
B.5 Format Error Analysis
Table 11 complements the quantitative results in Table A by presenting representative error patterns across all CRAFT variants. Base models predominantly fail due to missing closing tags (60%–85% of errors) and response truncation (30%–40% for complex templates), while GRPO dramatically reduces errors, with CRAFT and CRAFT achieving perfect compliance on 7B models.
B.6 CRAFT Template Sensitivity
As observed in Table 3.3.2, CRAFT shows limited but non-trivial gains at the smallest scale. Here we examine this model’s sensitivity to trace template choice in detail, revealing striking differences between SFT and GRPO failure modes. SFT training collapses on templates CRAFT, CRAFT, and CRAFT (EM 1.5% across benchmarks), achieving meaningful performance only on CRAFT (1.10/12.80/22.45%) and CRAFT (0.50/11.75/23.15%). In contrast, GRPO rescues templates CRAFT–CRAFT but still fails on CRAFT. Specifically, GRPO achieves 7.25/24.80/22.85% on CRAFT, 6.50/20.60/22.65% on CRAFT, 4.75/20.20/20.50% on CRAFT, and 8.45/25.95/24.70% on CRAFT, while CRAFT remains completely broken (0.00/0.00/0.05%).
This dichotomy reveals three key insights: (1) SFT cold-start is fragile for 0.5B models: templates with explicit <gold_docs> fields overwhelm supervised imitation, causing training collapse. (2) GRPO’s exploration mechanism overcomes SFT’s brittleness: reinforcement learning enables the model to discover valid reasoning paths even on complex structures, as long as some structural scaffolding exists. (3) CRAFT’s complete failure across both training paradigms confirms that 0.5B models fundamentally require explicit reasoning structure for stable learning. The optimal template for 0.5B deployment is CRAFT (<reason>+<answer>), providing the best balance of structure and simplicity.
| Model | Judge | MuSiQue | HotpotQA | 2WikiMHQA | |||
| Faith. | Faith. | Faith. | |||||
| Qwen2.5-7B | Qwen3-Max | 58.25 | – | 83.65 | – | 76.85 | – |
| Gemini-Flash | 55.40 | -2.85 | 82.15 | -1.50 | 73.90 | -2.95 | |
| Qwen2.5-7B | Qwen3-Max | 50.85 | – | 83.50 | – | 78.95 | – |
| Gemini-Flash | 49.70 | -1.15 | 81.40 | -2.10 | 77.50 | -1.45 | |
| CRAFT w/o judge | Qwen3-Max | 71.20 | – | 91.80 | – | 89.25 | – |
| Gemini-Flash | 69.85 | -1.35 | 90.25 | -1.55 | 86.80 | -2.45 | |
| CRAFT (full) | Qwen3-Max | 80.35 | – | 95.00 | – | 94.80 | – |
| Gemini-Flash | 77.60 | -2.75 | 93.60 | -1.40 | 93.25 | -1.55 | |
| Metric | MuSiQue | HotpotQA | 2WikiMHQA |
| Pearson | 0.998 | 0.999 | 0.997 |
| Spearman | 1.000 | 1.000 | 1.000 |
| Mean | 2.03 | 1.64 | 2.10 |
| Ranking Agreement | 4/4 | 4/4 | 4/4 |
Appendix C Cross-Judge Evaluation
To address the potential circularity of using Qwen3-Max as both the training reward provider and evaluation judge, we re-evaluate all model configurations using Gemini-2.5-Flash Google DeepMind (2025) as an independent judge with identical evaluation prompts. Gemini-2.5-Flash is not involved in any stage of CRAFT training.
Table 12 reports the faithfulness scores from both judges across all model configurations under CRAFT. Gemini-2.5-Flash assigns systematically lower absolute scores than Qwen3-Max (mean = 1.92), which is expected given known differences in evaluation stringency across models. Crucially, the relative ordering of all four configurations is fully consistent under both judges across all three benchmarks (Spearman = 1.0).
Table 13 confirms strong agreement: Pearson and Spearman across all benchmarks, demonstrating that the reported faithfulness improvements are robust to judge selection.
Appendix D Case Studies
| Aspect | Base Model (EM=0, Faith=0) | CRAFT (EM=1, Faith=1) |
[…]
Subq1: Where is the Oh Yeah album from?
Subq1: Where is the Oh Yeah performer from?
Subq2: What is the largest city of that state? Subq3: Who won the Indy car race there? |
[…]
| 1 - Plan steps followed in order | 1 - Plan followed | |
| 0 - Cites doc [1] not in gold_docs [5,12,14] | 1 - All cited docs in gold_docs | |
| 1 - Answer matches reasoning conclusion | 1 - Mario Andretti supported | |
| 0 - “Ryan Hunter-Reay won” not in doc [1] | 1 - All claims verified | |
| 0.50 | 1.0 | |
We present an additional case from MuSiQue comparing Qwen2.5-7B Base and CRAFT outputs. This case complements the right-answer-wrong-reason example in §4.6 by showing a case where both accuracy and faithfulness are corrected.
D.1 Case: Base (EM=0, Faith=0) CRAFT (EM=1, Faith=1)
Gold Answer: Mario Andretti